OpenStudy (anonymous):

Which set of rectangular coordinates describes the same location as the polar coordinates

OpenStudy (anonymous):

\[(3\sqrt{2},5\pi/4)?\]

OpenStudy (anonymous):

A. (-3,-3) B. (-3,3) C. (3,-3) D.

OpenStudy (anonymous):

d is \[(-3\sqrt{2},3\sqrt{2})\]

OpenStudy (anonymous):

A the answer